# Hypergraph adjusted plus-minus

### v1

@@ 5.3 Case Study
Player Name Minutes PM (rank) HAPM (rank) LAPM (rank)
Joel Embiid 2295 365 (1) 89.1 (1) 72.2 (1)
James Harden 793 151 (4) 43.3 (2) 58.6 (4)
Tyrese Maxey 2648 192 (2) 30.0 (3) 61.4 (2)
Tobias Harris 2542 170 (3) 27.1 (4) 60.7 (3)
Georges Niang 1742 69 (8) 0.1 (5) 56.4 (8)
@@
Table 5: Top five Sixers players in 2021-22 according to HAPM.
@@
We show the top five individual players according to HAPM in Table 5. Unsurprisingly, we see the top four players are Embiid, Harden, Maxey, and Tobias Harris, a former first-round draft pick. Importantly, despite only having a PM of 151, Harden is identified as the second best player by HAPM. Although there are positive rank correlations between our metrics and PM (0.65 and 0.93 for HAPM and LAPM, respectively), by plotting them against PM, we can identify players that may be over or under-valued based on PM alone. We plot this in Figure 4.
@@
Figure 4: HAPM (left) and LAPM (right) for individual players on the Sixers in 2021-22.
@@
The top six pairs according to HAPM are recorded in Table 6 and visualized in Figure 5 (the rank correlations with PM are 0.58 for both metrics). Here we see that not only is Embiid a great individual player, but he contributes to the top four best pairings (this is true for both HAPM, as shown, and LAPM). This is consistent with Embiid’s reputation as a player with the ability to pass out of a double team. Furthermore, despite only playing together for 600 total minutes, HAPM and LAPM recognize the successful pairing of Embiid and Harden.
@@
Interestingly, we also see a high ranking for the pairing of Embiid and Matisse Thybulle (8th best pair according to HAPM and 5th best pair according to LAPM) despite having a PM of only 82. This can be explained by the excellent defensive pairing of Embiid and Thybulle, which both of our methods are able to detect. Not surprisingly, the worst ranking pair according to HAPM (shown in Figure 4) is Deandre Jordan and Paul Reed, who are both big men with overlapping skillsets. It is also interesting that the two best pairs according to PM (Embiid and Maxey & Embiid and Harris) have relatively low pair rankings (10 and 8 respectively) according to LAPM. This may suggest that the high PM of these pairings can be attributed to individual player success or the success of higher-order generalized lineups (e.g. Embiid and Harris play together in the 3rd, 9th, 13th and 15th best ranked trios according to LAPM).
Player Pair PM (rank) HAPM (rank) LAPM (rank)
Joel Embiid:James Harden 193 (3) 26.5 (1) 57.6 (1)
Joel Embiid:Tyrese Maxey 268 (1) 25.3 (2) 56.4 (10)
Joel Embiid:Tobias Harris 256 (2) 25.2 (3) 56.5 (8)
Danny Green:Joel Embiid 171 (4) 19.6 (4) 57.5 (2)
Tyrese Maxey:James Harden 144 (8) 19.3 (5) 56.8 (6)
Tobias Harris:James Harden 154 (6) 19.1 (6) 57.0 (4)
@@
Table 6: Top six Sixers pairs in 2021-22 according to HAPM.
@@
Figure 5: HAPM (left) and LAPM (right) for pairs on the Sixers in 2021-22.
@@
Figure 6: Line graph visualization of player/pair (left) and player/trio (right) HAPM rankings for PHI 2021-2022. The size of the nodes corresponds to ranking within the individuals/pairs/trios. The blue nodes represent players, and the red nodes, labeled with rank, represent the pair/trio.
An immediate takeaway is the central role of Embiid, as he is part of 12 of the best 15 pairs and 14 of the 15 best trios according to HAPM. In both visualizations, he is at the center of the network. Exploratory analyses for other teams and other years reveals that this is very common; the superstars are the most central nodes in the visualizations. Further investigation reveals that pair nodes that are larger than one or both of the player nodes (Embiid and Danny Green; Embiid and Thybulle; Embiid and Seth Curry; Embiid and Shake Milton) may represent possible synergies, whereas pair nodes that are smaller than the player nodes (Embidd and Isaiah Joe; Harris and Maxey) might indicate inefficient pairings.
@@ 6 Discussion
In this paper, we propose a new framework for ranking individual players using box-score plus-minus data, while simultaneously ranking all genearlized lineups. We show that our methods, HAPM and LAPM, are more correlated with advanced metrics for ranking individual players than ordinary APM. We also provide a case study on the Philadelphia 76ers in 2021-22 that seemingly identifies the best individuals, pairs, and trios. We also show how our methods can easily be visualized in interpretable lineup networks.
There a few important things to note about our methods. First, as previously mentioned, PM, APM, and its variants including our proposed HAPM and LAPM are highly correlated with a team’s winning percentage. This is not inherently bad since winning is important, but this is also not surprising because these methods only use plus-minus data, whereas advanced metrics use many covariates. HAPM and LAPM provide an improvement over APM with respect to individual rankings, as measured by correlation to advanced metrics, while only using plus-minus data. For individuals, there is also the question of whether to rank using the model coefficients or the model predictions. In Table 7, we show that the coefficients $\widehat{\beta}_{\text{HAPM}}$ are strongly correlated (0.73) with the predicted values $\widehat{y}_{\text{HAPM}}$ for individuals. This suggests that there may be little difference in using one versus the other. This decision does not have to be made when utilizing LAPM, as the coefficients and predicted values are the same.
Figure 7: $\widehat{\beta}_{\text{HAPM}}$ vs the predicted values $\widehat{y}_{\text{HAPM}}$ for players on PHI 2021-2022.
@@
There are several directions for future research. Here, we presented two new metrics, LAPM and HAPM, for ranking generalized lineups. While we compared their results and stated the theoretical differences in the two approaches, we did not attempt to understand the cause of the metrics (dis)agreement. This may be a case-by-case exercise that could shed light on the inner workings of any one team. Furthermore, we did not discuss the uncertainty estimates in our model coefficients or fitted values. There could be a lot more information to uncover by examining the posterior distributions of the coefficients in LAPM. For example, one may be able to identify generalized lineups that are significantly better or worse than others, in a statistical sense, as opposed to providing a ranking. Furthermore, while it is straightforward to extend HAPM and LAPM to other APM variants that account for other non box-score statistics [17], it is less obvious how to incorporate defense for HAPM and LAPM for a league analysis rather than at the team level.

### v2

@@ 5.3 Case Study
Player Name Minutes PM HAPM (Rank CI) LAPM, Rank (Rank CI)
Joel Embiid 2295 365 14.1 (1 - 1) 166.7, 1 (1 - 1)
James Harden 793 151 8.9 (2 - 3) 73.9, 4 (4 - 5)
Tyrese Maxey 2648 192 8.0 (2 - 4) 94.2, 2 (2 - 3)
Tobias Harris 2542 170 7.8 (3 - 4) 88.6, 3 (3 - 3)
Danny Green 1349 77 3.0 (5 - 6) 56.8, 7 (6 - 10)
Seth Curry 1565 112 2.4 (5- 7) 68.3, 5 (4 - 6)
Matisse Thybulle 1685 88 1.9 (6 - 8) 60.1, 6 (6 - 8)
Georges Niang 1742 69 1.7 (7 - 8) 55.4, 8 (7 - 9)
@@
Table 5: Top eight Sixers players in 2021-22 according to HAPM (and 50% confidence intervals for implied rank).
@@
We show the top eight individual players according to HAPM in Table 5. We provide their HAPM and LAPM along with the 50% confidence intervals for the implied rank. The full distribution of HAPM/LAPM values is given in Figure A1 in the appendix. Unsurprisingly, we see the top four players are Embiid, Harden, Maxey, and Tobias Harris, a former first-round draft pick. Importantly, despite only having a PM of 151, Harden is identified as the second best player by HAPM. Although there are positive rank correlations between our metrics and PM (0.98 and 0.94 for HAPM and LAPM, respectively), by plotting them against PM in Figure 4, we can identify players that may be over or under-valued based on PM alone.
@@
Figure 4: HAPM (left) and LAPM (right) for individual players on the Sixers in 2021-22. Error bars on the y-axes represent inner-quartile range confidence intervals.
@@
We record the top ten pairs according to HAPM in Table 6 and visualize them in Figure 5. The rank correlations with PM are 0.58 and 0.59 for HAPM and LAPM, respectively. Here we see that not only is Embiid a great individual player, but he contributes to the top four and seven of the top ten best pairings according to HAPM (and the top three and seven of the top ten best pairings according to LAPM). This is consistent with Embiid’s reputation as a player with the ability to pass out of a double team. Furthermore, despite only playing together for 600 total minutes, HAPM and LAPM recognize the successful pairing of Embiid and Harden. As we hoped, our metrics identify Harden as a good teammate, belonging to three of the top six pairs for both HAPM and LAPM.
@@
Interestingly, we see a high ranking for the pairing of Embiid and Matisse Thybulle (9th best pair according to HAPM and 5th best pair according to LAPM) despite having a PM of only 82. This can be explained by the excellent defensive pairing of Embiid and Thybulle, which both of our methods are able to detect. Not surprisingly, the worst ranking pair according to HAPM (shown in Figure 4) is Deandre Jordan and Paul Reed, who are both big men with overlapping skillsets. It is also interesting that the two best pairs according to PM (Embiid and Maxey & Embiid and Harris) have relatively low pair rankings (10 and 8 respectively) according to LAPM. This may suggest that the high PM of these pairings can be attributed to individual player success or the success of higher-order generalized lineups (e.g. Embiid and Harris play together in the 3rd, 9th, 13th and 15th best ranked trios according to LAPM).
Player Pair PM HAPM (Rank CI) LAPM, Rank, (Rank CI)
James Harden:Joel Embiid 193 22.9 (1 - 2) 67.1, 1 (2 - 9)
Joel Embiid:Tyrese Maxey 268 22.0 (1 - 3) 54.5, 10 (20 - 26)
Joel Embiid:Tobias Harris 256 21.9 (2 - 3) 55.6, 8 (17 - 23)
Danny Green:Joel Embiid 171 17.0 (5 - 8) 65.9, 2 (2 - 9)
James Harden:Tyrese Maxey 144 16.8 (4 - 10) 59.4, 6 (6 - 24)
James Harden:Tobias Harris 154 16.7 (4 - 11) 61.2, 4 (4 - 20)
Joel Embiid:Seth Curry 131 16.4 (6 - 9) 64.5, 3 (3 - 10)
Joel Embiid:Matisse Thybulle 82 15.9 (6 - 10) 59.8, 5 (6 - 18)
Tobias Harris:Tyrese Maxey 166 15.8 (5 - 13) 52.7, 17 (31 - 38)
Georges Niang:Joel Embiid 133 15.7 (7 - 10) 53.5, 12 (26 - 33)
@@
Table 6: Top ten Sixers pairs in 2021-22 according to HAPM (and 50% confidence intervals for implied rank).
@@
Figure 5: HAPM (left) and LAPM (right) for pairs on the Sixers in 2021-22. Error bars on the y-axes represent inner-quartile range confidence intervals.
@@
Figure 6: Line graph visualization of player/pair (left) and player/trio (right) HAPM rankings for PHI 2021-22. The size of the nodes corresponds to ranking within the individuals/pairs/trios. The blue nodes represent players, and the red nodes, labeled with rank, represent the pair/trio.
An immediate takeaway is the central role of Embiid, as he is part of 12 of the best 15 pairs and 14 of the 15 best trios according to HAPM. In both visualizations, he is at the center of the network. Exploratory analyses for other teams and other years reveals that this is very common; the superstars are the most central nodes in the visualizations (also shown for the Boston Celtics in Figure A3 in the appendix). Further investigation reveals that pair nodes that are larger than one or both of the player nodes (Embiid and Danny Green; Embiid and Thybulle; Embiid and Seth Curry; Embiid and Shake Milton) may represent possible synergies, whereas pair nodes that are smaller than the player nodes (Embidd and Isaiah Joe; Harris and Maxey) might indicate inefficient pairings.
@@ 6 Discussion
In this paper, we propose a new framework for ranking individual players using box-score plus-minus data, while simultaneously ranking all genearlized lineups. We show that our methods, HAPM and LAPM, are similarly or more correlated with advanced metrics for ranking individual players than ordinary APM. We also provide a case study on the Philadelphia 76ers in 2021-22 that seemingly identifies the best individuals, pairs, and trios. Lastly, we illustrate how our methods can easily be visualized in interpretable lineup networks.
There a few important things to note about our methods. First, as previously mentioned, PM, APM, and its variants including our proposed HAPM and LAPM are highly correlated with a team’s winning percentage. This is not inherently bad since winning is important, but this is also not surprising because these methods only use plus-minus data, whereas advanced metrics use many covariates. HAPM and LAPM provide an improvement over APM with respect to individual rankings, as measured by correlation to advanced metrics, while only using plus-minus data.
@@
There are several directions for future research. Here, we presented two new metrics, LAPM and HAPM, for ranking generalized lineups. While we compared their results and stated the theoretical differences in the two approaches, we did not attempt to understand the cause of the metrics (dis)agreement. This may be a case-by-case exercise that could shed light on the inner workings of any one team. Furthermore, while it is straightforward to extend HAPM and LAPM to other APM variants that account for other non box-score statistics (Kubatko et al., 2007), it is less obvious how to incorporate defense for HAPM and LAPM for a league analysis rather than at the team level.
Appendix A Appendix
In the LAPM model specification, we have
$\beta$ $\sim N(0,~{}\lambda^{-1}L_{w}^{-1}).$
To evaluate the uncertainty in our modeling process, one option is to fix a value for $\sigma^{2}$, i.e. view it as a known value. Alternatively, we could extend the model definition by allowing $\sigma^{2}$ to be a random variable with an inverse scaled chi-squared prior:
$\beta$ $\sim N(0,~{}\lambda^{-1}L_{w}^{-1})$
$\frac{1}{\sigma^{2}}$ $\sim\Gamma(\nu/2,\frac{\nu\tau^{2}}{2}).$
For the analysis presented in the paper, we chose to use a non-informative prior for $\sigma^{2}$, setting $\nu=\tau^{2}=0.$ We utilized the Metropolis adjusted Langevin algorithm (Girolami and Calderhead 2011) to sample from the posterior distributions. The posterior distributions displayed throughout the manuscript are based on 1000 samples after disregarding 10% for a burn-in period and thinning every 5 samples.
We note that while the center and spread of the posterior distributions (and therefore, the fitted values) are sensitive to prior specification, the resulting rankings of the generalized lineups are not. Also, while we set the hyper-prior parameters to be the same for all teams, it may be worthwhile to investigate the parameters at the team level if that is where one’s interests primarily lie.
Appendix B Appendix
Figure A1: HAPM (left) and LAPM (right) for individuals on the Sixers in 2021-22. The boxplots represent the distribution of HAPM and LAPM values obtained via bootstrap and MCMC, respectively.
Figure A2: HAPM (left) and LAPM (right) for the top pairs on the Sixers in 2021-22. The boxplots represent the distribution of HAPM and LAPM values obtained via bootstrap and MCMC, respectively.
Figure A3: Line graph visualization of player/pair (left) and player/trio (right) HAPM rankings for BOS 2021-22. The size of the nodes corresponds to ranking within the individuals/pairs/trios. The green nodes represent players, and the white nodes, labeled with rank, represent the pair/trio.